Solve: x+2(3x-1)=6x-5
Answer by rapaljer(4671)

x+2(3x-1)=6x-5
x+6x-2=6x-5
7x-2=6x-5

Subtract 6x from each side:
7x-6x-2=6x-6x-5
x-2=-5

Add +2 to each side:
x-2+2=-5+2
x=-3

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus